Suppose that I flip a fair coin 21 times. What is the probability that it will land on heads exactly 12 times?

Round to 4 decimal places.
Answer #1

SOLUTION

Given : fair coin = p(head) = q (tail) = 1-p=

number of coin tossed (n) = 21

formula : P(X = z) = C, * p* *q- where, Cy=- n! c!(n-1)!

find : Peractly 12) = P(X = 12) =?

P(X = 12) - C»* () ** (?)

(121-12 21! P(X = 12) = 1 12!(21 – 12)!

P(X – 12) – 12.47*()*()

21 * 20 * 19 * 18 * 17 * 16 * 15 * 14 * 13 * 12! P(X = 12) = ? 12! * (9*8*7*6*5*4*3*2*1)

P(X = 12) = 293930*

P(X = 12) = 0.1402

The probability that it will land on heads exactly 12 times is 0.1402